# HOLISMOKES XVI: Lens search in HSC-PDR3 with a neural network committee and post-processing for false-positive removal

### v1

@@ 3.2 Cleaning using modeling CNN
In parallel, we applied the ResNet from S23a to all 22,393 network candidates. This modeling network is trained on realistic mock images created in a similar way as those used for training the committee network, and obtained great performance in measuring the Einstein radii of lens systems when compared to gleeauto.py (Schuldt et al., 2023b), a code that relies on the well-tested lens modeling software Glee (Suyu & Halkola, 2010; Suyu et al., 2012a). The modeling network predicts the lens mass center and , the lens mass ellipticity and , and Einstein radius of a SIE profile as well as an external shear ( and ), and the corresponding 1 uncertainties.
Since the modeling network is trained on solely lens images (i.e. not in combination with lens classification as in Andika et al. (2024)), it is forced to provide reasonable model parameters (e.g., [05,) even if the given image is clearly not a lens. However, for such a given non-lens image, the modeling network is uncertain and the predicted errors can be significantly higher. Consequently, we can use this fact to reject false-positives from the classification network. To be conservative, we define the cuts here such that we do not exclude any lens candidate from our test set as shown in Fig. 4, which results in an additional rejection of 2,794 systems. We show some examples of rejected candidates in the bottom row of Fig. 3.
In sum, we reject around 29% of the network candidates as false positives by keeping the same TPR on our test sample. This cleaning is performed through automated and fast scripts and thus are also scalable for sample sizes expected from ongoing and upcoming wide-field imaging surveys which are two orders of magnitude higher. While excluding of the contaminants by keeping the same TPR is already a good improvement in the purity, we note that for significantly higher samples more stringent cuts can be applied to reject more contaminants with the downside of possibly excluding some lenses.
@@ 4 Visual inspection
Before the visual inspection of the remaining network candidates from Sect. 3, we excluded duplicates among the network candidates. Although CNNs are known to be translation invariant, we only excluded duplicates within two pixels when creating the 110 million parent sample to be very conservative. This ensures that we will not miss any candidate from the network due to a lower score if the lens is not well centered or partly out of the cutout. However, for the visual inspection, we removed duplicate cutouts within 2″. In order to preserve lines of sight with the most extended, and most likely deflector galaxy in the center, we rank ordered the catalog entries by decreasing -Kron radius and removed duplicates of lower rank. Excluding these duplicates reduces the human inspection time further, while the moderate cut of 2″ensures that we will not shift any multiple image out of the inspected cutout that would risk missing the identification.
Furthermore, we excluded additional image stamps previously inspected as part of our HOLISMOKES strong lens searches in earlier HSC data releases (C21; S22; Schuldt et al. (2025), hereafter S25) to lower the need in human resources, as we grade the systems in a similar way and would therefore have similar resulting grades. We only kept 50 lenses (high-quality grade A or B lens candidates, which we simply refer to as “lenses” hereafter) and 50 non-lenses from C21 and S25 relying both on the same network, as well as 50 lenses and 50 non-lenses from S22 for comparison222we use here a lower limit of the threshold score to include also a few systems that would be usually excluded with the cut at .. This resulted in a catalog of 11,874 candidates for visual inspection.
The visual inspection follows closely the procedure proposed by S25 and was conducted jointly with candidates presented by Andika et al. (2024). In short, we first carried out a calibration round using 200 system inspected by all eight inspectors333The visual inspectors are in alphabetical order by last name: I. T. A., S. B., R. C., C. G., A. M., S. S., S. H. S., and S. T. since we introduced several new features in the grading tool as detailed below. As previously, we use a three-panel image to show two different scalings of the bands, and, in contrast to previous inspections, an image of the bands to ease the identification of lensed quasars. While these image stamps have a size of , we additionally show stamps in the same filters and scaling as in the smaller cutout.
@@ 4.1 Binary inspection
We then split the catalog among ourselves, such that each system is inspected by two graders in a binary classification to exclude rapidly the majority of obvious non-lenses. As mentioned above, the catalog contains 200 systems from our previous searches, with 100 lenses and 100 non-lenses. Of the 100 lenses (or lens candidates), we recovered 98/100 with previous grade A or B and only missed two systems, namely, systems HSC J222002+060506444Listed as HSC J2220+0605 in C21. and HSC J090929+010030 which previously obtained both an average grade of 1.6 while the lower limit for grade B is 1.5. Of the 100 non-lenses, we forwarded 59/100, which, however, is understandable as several of these galaxies obtained previously an average grade slightly below 1.5 and we aimed on forwarded now all systems above a grade of 1 in the final grading scheme. Interestingly, we also forwarded 12/100 with previously average grade 0. Since this first stage is graded conservatively to rule out obvious non-lenses and there is subsequently another round of inspection for refinement, this result on the 200 systems is overall very good.
@@ 4.2 Multi-class inspection
In the next round, four individuals inspected the remaining 3,408 systems from the binary classification. At this stage, each inspector provided one of four possible grades ( corresponding to “no lens”, to “possible lens”, to “probable lens”, and to “definite lens”) and voted in case the lens is significantly offset from the cutout center. In total, 309 systems obtained the “off-center” flag by at least one person, such that the lens center was subsequently corrected. A re-centering at this stage was crucial for the subsequent round of grading by the other four remaining graders, as we also showed in the central image the lens center and Einstein radius predicted by the ResNet from S23a (see bottom row of Fig. 3 for examples).
Based on the average grade obtained from these four grades, we forwarded all systems with either or but standard deviation of the four grades above 0.75. Including those with lower grade but high discrepancy increases the forwarded sample significantly, but minimises the possibility of missing any good lens candidate. This resulted in 686 candidates, including 160 systems that got shifted, which were inspected by the remaining four graders. In addition to the grading and “off-center” flag from the previous round, these four individuals were tasked to also indicate if the predicted lens center or Einstein radius is significantly mis-predicted. Furthermore, for each candidate, two histograms of photometric redshifts from the lens candidate environment, one up to and the other up to , were shown in order to obtain a classification of the environment as well. As examples, the histograms of one system in an overdensity (bottom row) and one in the field (top row), are shown in Fig. 5. This allows us to assess if the predicted mass model, which we provide in Table 2 (see also Table 3 for explanations), is reliable and to characterise their environment, while increasing the visual inspection time only very little.
@@
Column | Header | Description |
(1) | Name | name for grade A and B lens candidates |
(2) | RA [deg] | Right Ascension in degrees of the inspected candidate |
(3) | Dec [deg] | Declination in degrees of the inspected candidate |
(4) | p | average score predicted by the network committee |
(5) | G | average grade between 0 and 3 obtained though visual inspection (see Sect. 4) |
(6) | sigma_G | standard deviation of obtained visual inspection grades. |
(7) | N_graders | number of visual inspectors (see Sect. 4) |
visual inspection grade from our earlier work S25 | ||
visual inspection grade from our earlier work Shu et al. (2022) | ||
visual inspection grade from our earlier work Cañameras et al. (2021) | ||
(17) | x_err | 1- value for the -center coordinate predicted by the modeling network from S23a |
(19) | y_err | 1- value for the -center coordinate predicted by the modeling network from S23a |
(21) | ex_err | 1- value for the -component of the complex ellipticity predicted by the modeling network from S23a |
(23) | ey_err | 1- value for the -component of the complex ellipticity predicted by the modeling network from S23a |
(25) | rE_err | 1- value of the Einstein radius predicted by the modeling network from S23a |
(27) | gma1_err | 1- value of the -component predicted by the modeling network from S23a |
(29) | gam2_err | 1- value of the -component predicted by the modeling network from S23a |
(31) | z | photometric redshift value from the catalog compiled catalog by S25 |
based on DEmP (Hsieh & Yee, 2014), Mizuki (Tanaka et al., 2018), and NetZ (Schuldt et al., 2021) | ||
(36) | Ntot | Sum of objects with photo- within a box of 200″on a side. |
http://sunguoyou.lamost.org/glc.html.). |
@@ 4.3 Comparison to previous inspected candidates
As mentioned above, we included 200 systems that we inspected already in previous works. From the 100 candidates that we previously excluded (i.e. that obtained an average grade ), only one system obtained now an average grade above 1.5 and is consequently listed as grade B candidate. From the sample of previously classified grade A and B lenses, we notice a tendency towards lower grades consistently among most graders. While there were some changes compared to previous inspections, such as moving from PDR2 to PDR3 and showing two color images and one color image instead of three color images, we ruled these aspects out by directly comparing some inspected images from S25 and this work. We note also that the applied scalings depend on the pixel values of the given cutout, such that a shifted image appears slightly different. Additionally, in S22 we focused on high- lenses and applied slightly different scaling functions, but find this to be not a major reason. Furthermore, we speculate that this tendency may come from the relatively pure sample (i.e., a higher proportion of lenses) that we had after the binary inspection, but note that we also had in previous works a binary inspection cleaning carried out by a single inspector, resulting in a comparably pure sample. In contrast, we noticed that the 80″80″cutouts help in specific cases to reject non-lenses as the overall environment is visible. Finally, we remark that it is known (Rojas et al., 2023; Schuldt et al., 2025) that every grader has a different expectation and also the grades among a single inspector varies. Consequently, it might be simply that our expectations for an object to qualify as a lens candidate has slightly increased over time, possibly because of the increasing sample of lenses and the higher availability of high-resolution images (e.g., Euclid, see Acevedo Barroso et al. (2024); Nagam et al. (2025), Walmsley et al. in prep.).
@@ 4.4 Final lens compilation
Despite the tendency towards lower grades, we adopt our traditional cuts to define grade A () and grade B (), and release in Table 2 the full catalog of visually inspected systems. This catalog includes jointly inspected systems that obtained a committee score and were also detected by VariLens (Andika et al., 2024, noted in column 12 of the catalog), but we exclude any duplicates within ″. Consequently, using the average grades of all eight inspectors to obtain a stable average (see also Rojas et al., 2023; Schuldt et al., 2025), we found 42 grade A and 187 grade B lens candidates, which are listed in Table 2. Furthermore, the newly discovered grade A candidates are shown in Fig. 6, while the grade B systems are shown in Fig. 7. We further provide in this table all grade A and B lens candidates from C21 (marked in column 15, see Table 3), S22 (marked in column 14), and S25 (marked in column 13) which we detected with our network committee as candidate but excluded for visual inspection (see also Sect. 4.2). In sum, we found with our network committee 95 grade A and 503 grade B lenses. From these systems, 506 systems (included not re-graded systems) were already known, while we show in Figs. 8 (grade A) and 9 (grade B) those that we re-graded in our inspection (either because they are in the small comparison sample, see Sect. 4.3, or because they were discovered by other work and not by C21, S22, or S25). This high recovery rate is expected, given the enormous lens search projects that already exploited HSC data and in particular our previous searches using individual networks that entered into our network committee now. However, we remind the reader that the new lens identification is only one aspect of this work.
@@ 4.5 Discussion on the lens properties
We further classified the lens environment using photo- histograms during the inspection by the second team (see Fig. 5 for examples), leading to an identification of 147 candidates in an overdensity (corresponding to at least two votes) of which 22 and 66 are ranked as grade A and B candidates, respectively. With the automated photo- procedure to identify overdensities presented by S25, we identify 569 systems among the whole inspected sample, of which 18 and 70 are classified as grade A or B lens candidates, respectively (see also Tab. 1). While the automated procedure can be easily applied to large samples, the visual classification requires in particular human time that is highly limited. On the other hand, we visually identified several systems to be in an overdensity that are missed by the automated procedure due to missing or band observations required for the photo- codes. In total, we find a good overlap between the two complementary procedures.
As mentioned earlier, we further showed the network predicted lens center and Einstein radius during final inspection. Since we found in the calibration round that the model network from Schuldt et al. (2023a) has difficulties to correctly predict the model if the lens is not well centered, we mitigated this through a re-centering between the last two inspection stages. We overlay the predicted lens center and Einstein radius on the cutouts of our lens candidates in Appendix B. For the reported lens center coordinates in Table 2, we only found 9% of the lens candidates to have a relatively poor predicted lens center or Einstein radius which might be because of a group- or cluster-scale lens not suited for the modeling network or a remaining offset of the lens. We provide the predicted values of the seven mass model parameters with their corresponding 1- uncertainties in Table 2 (columns 16 to 29, see also Table 3) as well as a flag (column 30) if the lens center and Einstein radius are matching reasonably the system (defined as two or more votes). This demonstrates once more the power of the modeling network and shows that the result can be used for further analysis.
While we applied in Sect. 3.2 relatively weak cuts on the model parameter uncertainties (see Fig. 4), we note that more stringent cuts are possible when the sample sizes are becoming significantly larger. This would lead to an even higher fraction of systems that can be easily ruled out, while keeping most of the lens candidates.
@@ 5 Summary and Conclusion
While the resulting sample size would have been easily possible to visually inspect, this will change with the next generation of wide-field imaging surveys, such that further cleaning is unavoidable. We tested two approaches, first using SExtractor and the HSC pixel-level flags to reject images with mostly non-zero background or image artefacts, and second, the ResNet modeling network from Schuldt et al. (2023a) as it predicts higher uncertainties for non-lenses. This lowered the contaminants by 29%, while not excluding any lens candidate from our previous identifications. Such post-processing scripts are expected to play a crucial role when two orders of magnitude more images need to be classified.
Thanks to a visual inspection of the cleaned network candidate list, we identified 95 grade A and 503 grade B candidates, which include 506 previously known systems. We provide in Table 2 their coordinates together with their SIE and external shear parameters obtained from the ResNet presented by Schuldt et al. (2023a). During visual inspection, we also showed, which was done for the first time, the predicted lens center and Einstein radius and characterized the reliability. We found that only % of the candidates obtained a lens center or Einstein radius that is not well predicted, once the system is well centered. Moreover, we characterized their environment with two complementary approaches, either through visual inspection of histograms showing the photo- values of their surrounding, or the selection cuts elaborated by S25. In both cases, we find 88 grade A or B systems to be in a significant overdense environment. By releasing Table 2, which includes all visually inspected candidates with our obtained average grades and their characteristics as well as their previous discoveries, we release one of the most complete catalogs of strong lensing systems so far.
Particularly in preparation for the ongoing and upcoming wide-field imaging surveys, the removal of false-positives of network classifiers will be a challenge, such that the approaches presented here are expected to play a crucial role. Furthermore, fast and autonomous techniques to characterize and analyze the new systems, as tested here as well, will be the challenge we face.
@@
Table 2 will be available after acceptance of the publication at the CDS via anonymous ftp to cdsarc.cds.unistra.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/. It will be further released through the HOLISMOKES collaboration webpage at www.holismokes.org and accordingly incorporated in the SuGOHI data base accessible through https://www-utap.phys.s.u-tokyo.ac.jp/~oguri/sugohi/.

### v2

@@ 3.2 Cleaning using modeling CNN
In parallel, we applied the ResNet from S23a to all 22,393 network candidates. This modeling network is trained on realistic mock images created in a similar way to the ones used for training the committee network, and obtained a great performance in measuring the Einstein radii of lens systems when compared to gleeauto.py (Schuldt et al., 2023b), a code that relies on the well-tested lens modeling software Glee (Suyu & Halkola, 2010; Suyu et al., 2012a). The modeling network predicts the lens mass center, and , the lens mass ellipticity, and , and Einstein radius, , of a SIE profile as well as an external shear ( and ), and the corresponding 1 uncertainties.
Since the modeling network is trained on solely lens images (i.e., not in combination with lens classification as in Andika et al. (2025)), it is forced to provide reasonable model parameters (e.g., [05, ) even if the given image is clearly not a lens. However, for such a given non-lens image, the modeling network is uncertain and the predicted errors can be significantly higher. Consequently, we can use this fact to reject false-positives from the classification network. To be conservative, we define the cuts here such that we do not exclude any lens candidate from our test set, as is shown in Fig. 4, which results in an additional rejection of 2,794 systems. We show some examples of rejected candidates in the bottom row of Fig. 3.
In sum, we rejected around 29% of the network candidates as false positives by keeping the same TPR on our test sample. This cleaning was performed through automated and fast scripts, and thus is also scalable for sample sizes expected from ongoing and upcoming wide-field imaging surveys that are two orders of magnitude higher. While excluding of the contaminants by keeping the same TPR is already a good improvement in the purity, we note that for significantly higher samples more stringent cuts can be applied to reject more contaminants, with the downside of possibly excluding some lenses.
@@ 4 Visual inspection
Before the visual inspection of the remaining network candidates from Sect. 3, we excluded duplicates among the network candidates. Although CNNs are known to be translation-invariant, we only excluded duplicates within two pixels when creating the 110 million parent sample to be very conservative. This ensured that we got network scores for all individual galaxies in the parent sample, and that we preserved cutouts centered on the actual lens galaxies. However, for the visual inspection, we removed duplicate cutouts within 2″. In order to preserve lines of sight with the most extended and most likely deflector galaxy in the center, we rank ordered the catalog entries by decreasing -Kron radius and removed duplicates of lower rank. Excluding these duplicates reduces the human inspection time further, while the moderate cut of 2″ensures that we will not shift any multiple image out of the inspected cutout, which would risk us missing the identification.
Furthermore, we excluded additional image stamps previously inspected as part of our HOLISMOKES strong lens searches in earlier HSC data releases (C21; S22; Schuldt et al. (2025), hereafter S25) to lower the need in human resources, as we grade the systems in a similar way and would therefore have similar resulting grades. We only kept 50 lenses (high-quality grade A or B lens candidates, which we simply refer to as “lenses” hereafter) and 50 non-lenses from C21 and S25 relying both on the same network, as well as 50 lenses and 50 non-lenses from S22 for comparison333We use here a lower limit of the threshold score to include also a few systems that would be usually excluded with the cut at .. This resulted in a catalog of 11,874 candidates, including the 200 systems mentioned above, for visual inspection.
The visual inspection closely followed the procedure proposed by S25 and was conducted jointly with candidates presented by Andika et al. (2025). In short, we (1) carried out a calibration round using 200 systems inspected by all eight inspectors444The visual inspectors are in alphabetical order by last name: I. T. A., S. B., R. C., C. G., A. M., S. S., S. H. S., and S. T. since we introduced several new features in the grading tool, as is detailed below. The actual grading then comprised (2) a binary classification (Sect. 4.1), a (3) four-grade inspection of four individuals, and (4) another four-grade inspection from the remaining inspectors, which additionally characterized the predicted mass model and the candidate environment (see Sect. 4.2). Finally, we obtained the average grade, , from the eight individual grades collected for the interesting candidates (reported in Table 2). As previously, we use a three-panel image to show two different scalings of the bands, and, in contrast to previous inspections, an image of the bands to ease the identification of lensed quasars. While these image stamps have a size of , we additionally show stamps in the same filters and scaling as in the smaller cutout.
@@ 4.1 Binary inspection
We then split the catalog among ourselves, such that each system was inspected by two graders in a binary classification to rapidly exclude the majority of obvious non-lenses. As it was mentioned above, the catalog contains 200 systems from our previous searches, with 100 lenses and 100 non-lenses. Of the 100 lenses (or lens candidates), we recovered 98/100 with a previous grade of A or B and only missed two systems, namely systems HSC J222002+060506555Listed as HSC J2220+0605 in C21. and HSC J090929+010030 which previously obtained both an average grade of 1.6, while the lower limit for grade B is 1.5. Of the 100 non-lenses, we forwarded 59/100, which, however, is understandable as several of these galaxies previously obtained an average grade slightly below 1.5 and we aim to now forwarded all systems above a grade of 1 in the final grading scheme. Interestingly, we also forwarded 12/100 that previously had an average grade of 0. Since this first stage is graded conservatively to rule out obvious non-lenses and there is subsequently another round of inspection for refinement, this result for the 200 systems is overall very good.
@@ 4.2 Multi-class inspection
In the next round, four individuals inspected the remaining 3,408 systems from the binary classification. At this stage, each inspector provided one of four possible grades ( corresponding to “no lens”, to “possible lens”, to “probable lens”, and to “definite lens”) and voted in case the lens was significantly offset from the cutout center. In total, 309 systems obtained the “off-center” flag by at least one person, such that the lens center was subsequently corrected. A re-centering at this stage was crucial for the subsequent round of grading by the other four remaining graders, as we also showed in the central image the lens center and Einstein radius predicted by the ResNet from S23a (see bottom row of Fig. 3 for examples).
Based on the average grade, , obtained from these four grades, we forwarded all systems with either or but a standard deviation of the four grades above 0.75. Including those with a lower grade but a high discrepancy increases the forwarded sample significantly, but minimizes the possibility of missing any good lens candidate. This results in 686 candidates, including 160 systems that got shifted, which were inspected by the remaining four graders. In addition to the grading and “off-center” flag from the previous round, these four individuals were tasked with also indicating if the predicted lens center or Einstein radius is significantly mis-predicted. Furthermore, for each candidate, two histograms of photometric redshifts from the lens candidate environment, one up to and the other up to , were shown in order to obtain a classification of the environment as well. As examples, the histograms of one system in an overdensity (bottom row) and one in the field (top row) are shown in Fig. 5. This allows us to assess if the predicted mass model, which we provide in Table 2 (see also Table 3 for explanations), is reliable and to characterize the systems’ environment, while only very slightly increasing the visual inspection time.
@@ 4.3 Comparison to previous inspected candidates
As was mentioned above, we included 200 systems that we already inspected in previous works. From the 100 candidates that we previously excluded (i.e., that obtained an average grade of ), only one system has now obtained an average grade above 1.5 and is consequently listed as a grade B candidate. From the sample of previously classified grade A and B lenses, we notice a tendency toward lower grades among most graders. While there were some changes compared to previous inspections, such as moving from PDR2 to PDR3 and showing two color images and one color image instead of three color images, we checked these aspects by directly comparing some inspected images from S25 and this work (see example images in Fig. 14), and do not see this as a major source of bias. We also note that the applied scalings depend on the pixel values of the given cutout, such that a shifted image appears slightly different. Additionally, in S22 we focused on high- lenses and applied slightly different scaling functions, but find this to be not a major reason. Furthermore, we speculate that this tendency may come from the relatively pure sample (i.e., a higher proportion of lenses) that we had after the binary inspection, but note that we also had in previous works a binary inspection cleaning carried out by a single inspector, resulting in a comparably pure sample. In contrast, we noticed that the 80″80″cutouts help in specific cases to reject non-lenses as the overall environment is visible. Finally, we remark that it is known (Rojas et al., 2023; Schuldt et al., 2025) that every grader has a different expectation and also the grades among a single inspector vary. Consequently, it might be simply that our expectations of an object qualifying as a lens candidate have slightly increased over time, possibly because of the increasing sample of lenses and the higher availability of high-resolution images (e.g., Euclid, see Acevedo Barroso et al. (2025); Nagam et al. (2025); Euclid Collaboration et al. (2025b)).
@@ 4.4 Final lens compilation
Despite the tendency toward lower grades, we adopt our traditional cuts to define grade A () and grade B (), and release in Table 2 the full catalog of visually inspected systems. This catalog includes jointly inspected systems that obtained a committee score of and were also detected by VariLens (Andika et al., 2025, noted in column 12 of the catalog), but we exclude any duplicates within ″. Consequently, using the average grades, , of all eight inspectors to obtain a stable average (see also Rojas et al., 2023; Schuldt et al., 2025), we found 42 grade A and 187 grade B lens candidates, which are listed in Table 2. Furthermore, the newly discovered grade A candidates are shown in Fig. 6, while the grade B systems are shown in Fig. 7. We further provide in this table all grade A and B lens candidates from C21 (marked in column 15, see Table 3), S22 (marked in column 14), and S25 (marked in column 13) that we detected with our network committee as candidates but excluded from visual inspection (see also Sect. 4.2). In sum, we found with our network committee 95 grade A and 503 grade B lenses. From these systems, 506 systems (included not regraded systems) were already known, while we show in Figs. 8 (grade A) and 9 (grade B) those that we regraded in our inspection (either because they are in the small comparison sample, see Sect. 4.3, or because they were discovered by other work and not by C21, S22, or S25). This high recovery rate is expected, given the enormous lens search projects that already exploited HSC data and in particular our previous searches using individual networks that entered into our network committee now. However, we remind the reader that the new lens identification is only one aspect of this work.
@@ 4.5 Discussion on the lens properties
We further classified the lens environment using photo- histograms during the inspection by the second team (see Fig. 5 for examples), leading to an identification of 147 candidates in an overdensity (corresponding to at least two votes), of which 22 and 66 are ranked as grade A and B candidates, respectively. With the automated photo- procedure to identify overdensities presented by S25, we identify 569 systems among the whole inspected sample, of which 18 and 70 are classified as grade A or B lens candidates, respectively (see also Tab. 1). While the automated procedure can be easily applied to large samples, the visual classification requires in particular human time that is highly limited. On the other hand, we visually identified several systems as being in an overdensity that were missed by the automated procedure due to missing or band observations required for the photo- codes. In total, we find a good overlap between the two complementary procedures.
As was mentioned earlier, we further showed the network-predicted lens center and Einstein radius during final inspection. Since we found in the calibration round that the model network from Schuldt et al. (2023a) has difficulties in correctly predicting the model if the lens is not well centered, we mitigated this through a re-centering between the last two inspection stages. We overlay the predicted lens center and Einstein radius on the cutouts of our lens candidates in Appendix C. For the reported lens center coordinates in Table 2, we only found 9% of the lens candidates to have a relatively poor predicted lens center or Einstein radius, which might be because of a group- or cluster-scale lens not being suited for the modeling network or a remaining offset of the lens. We provide the predicted values of the seven mass model parameters with their corresponding 1 uncertainties in Table 2 (columns 16 to 29, see also Table 3) as well as a flag (column 30) if the lens center and Einstein radius match the system reasonably (defined as two or more votes). This demonstrates once more the power of the modeling network and shows that the result can be used for further analysis.
While we applied in Sect. 3.2 relatively weak cuts on the model parameter uncertainties (see Fig. 4), we note that more stringent cuts are possible when the sample sizes become significantly larger. This would lead to an even higher fraction of systems that can be easily ruled out, while keeping most of the lens candidates.
@@ 5 Summary and conclusion
While it would easily have been possible to visually inspect the resulting sample size, this will change with the next generation of wide-field imaging surveys, such that further cleaning is unavoidable. We tested two approaches, first using SExtractor and the HSC pixel-level flags to reject images with mostly nonzero background or image artifacts, and second, the ResNet modeling network from Schuldt et al. (2023a) as it predicts higher uncertainties for non-lenses. This lowered the contaminants by 29%, while not excluding any lens candidate from our previous identifications. Such post-processing scripts are expected to play a crucial role when two orders of magnitude more images need to be classified.
Thanks to a visual inspection of the cleaned network candidate list, we identified 95 grade A and 503 grade B candidates, including 506 previously known systems. We provide in Table 2 their coordinates together with their SIE and external shear parameters obtained from the ResNet presented by Schuldt et al. (2023a). During the visual inspection, we also showed for the first time the predicted lens center and Einstein radius, and characterized the reliability. We found that only % of the candidates obtained a lens center or Einstein radius that is not well predicted, once the system is well centered. Moreover, we characterized their environment with two complementary approaches, either through a visual inspection of histograms showing the photo- values of their surroundings, or the selection cuts elaborated by S25. In both cases, we find 88 grade A or B systems to be in a significantly overdense environment. In Table 2, which includes all visually inspected candidates with our obtained average grades and their characteristics as well as their previous discoveries, we are releasing one of the most complete catalogs of strong lensing systems so far.
Particularly in preparation for the ongoing and upcoming wide-field imaging surveys, the removal of false-positives of network classifiers will be a challenge, so the approaches presented here are expected to play a crucial role. Developing and testing fast and autonomous techniques to characterize and analyze the new systems, such as the ones tested here as well, will be a necessity.
@@
Table 2 is available at the CDS via anonymous ftp to cdsarc.cds.unistra.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/. It was further released through the HOLISMOKES collaboration webpage at www.holismokes.org and accordingly incorporated in the SuGOHI data base accessible through https://www-utap.phys.s.u-tokyo.ac.jp/~oguri/sugohi/.